OpenStudy (anonymous):

Multiply x^1/2 times x^3/7?

OpenStudy (mertsj):

Use this property: \[a^m \times a^n=a ^{m+n}\]

OpenStudy (anonymous):

so I just add them?

OpenStudy (mertsj):

In other words, add the exponents.

OpenStudy (anonymous):

awesome! So is it x^4/9

OpenStudy (mertsj):

no

OpenStudy (mertsj):

\[\frac{1}{2}+ \frac{3}{7}=\]

OpenStudy (mertsj):

Get a common denominator and then add the numerators.

OpenStudy (anonymous):

so a common denominator would be 14 and you would get 7/14 x 6/14?

OpenStudy (anonymous):

so 13/14?

OpenStudy (mertsj):

7/14 PLUS 6/14=13/14

OpenStudy (anonymous):

oh, oops thats what I meant. Thank you so much!

OpenStudy (mertsj):

yw